OpenStudy (anonymous):

Side one of a triangle is (3x+10), side two of a triangle is (5x) and side 3 of a triangle is (2x + 25). If the perimeter of the triangle is 200, find the value of x. Answer 16.5 59.5 82.5 58

OpenStudy (anonymous):

(3x+10)+(5x)+(2x+25)=200 10x+35=200 10x=200-35 10x=165 x=16.5

OpenStudy (anonymous):

thnk you

OpenStudy (anonymous):

sure thing